Assume that the production function is given by Q = 3K + 4L. What is the average product
(“Average product”) of capital (K) when 10 units of capital and 10 units of work are used?
Ear: Select the correct average product formula. As always: Present the formula that
Let you solve the problem, all your calculations and write a final sentence explaining the
result. (5 points) 
Answer #1

8 = 3K+ 4L Average Product - 8 = 34+4 When K= L = 10 Average Product = 3x10 + 4 10 The of Tall total output produced labor is
Answer #2
Add you answer here... producaion f uncation
source: student
answered by: buze